OpenStudy (anonymous):

What is the 10th term in this sequence: A(n) = -5 + (n - 1)(2) whatt?

OpenStudy (anonymous):

put n=10 in abov equation A(10)=-5+(10-1)(2) =-5+18 =-13

OpenStudy (anonymous):

so the 10th term is a(10)=-13

OpenStudy (anonymous):

\(A_n=-5+(n-1)^2\) or \[A_n=-5+2(n-1)\]?